# How Hyper-Datafication Impacts the Sustainability Costs in Frontier AI

### v2

@@ 3.3. External Data Sources
We extract global annual investment in data centre infrastructure from the International Energy Agency (IEA) (International Energy Agency, 2025b). To characterise the geographic concentration and expansion of data centres, we use records from the Data Centre Map (19). We assessed electricity demand associated with data centres using historical and projected estimates from the IEA (International Energy Agency, 2024a) and focus on electricity consumption as a primary indicator of environmental impact associated with data centre expansion. We source the carbon intensity for electricity generation from Our World in Data (Ritchie et al., 2023a).
@@ 4.1. Environmental Costs
Figure 2. Left: Estimated provider-side storage energy (GWh). Right: Estimated user-side storage energy (TWh), assuming that 10 percent of downloads result in three months of local storage. Right axes show human-equivalent annual emissions. To provide a sense of scale, we relate cumulative emissions to the global average annual per-capita footprint of 4.73 tCO2eq (Ritchie et al., 2023b).
@@
Figure 2 reports estimated provider-side (Left) and user-side storage (Right) energy consumption associated with Hugging Face datasets. These estimations are based on the assumptions and calculations provided in Appendix D.2. Provider-side storage energy consumption remains modest in absolute terms but increases steadily as datasets accumulate, having reached approximately 1 GWh. User-side costs are approximately three orders of magnitude larger (>>2 TWh), driven by the repeated number of downloads across a large user base.
@@
Carbon intensity varies across electricity grids due to the differences in the energy mix. Storing all Hugging Face data in the EU rather than the US would reduce the associated carbon footprint by approximately 38%, due to the lower average carbon intensity in the EU (237 gCO2eq/kWh) compared with the US (384 gCO2eq/kWh) in 2024 (Ritchie et al., 2023a). Depending on the location of data centre infrastructure, storage-related carbon footprints can differ by up to a factor of 30 across regions [^7] This reflects contrasts between high- and low-carbon electricity grids, such as Kosovo (959 gCO2eq/kWh) and Norway (31 gCO2eq/kWh) in 2024 (Ritchie et al., 2023a).. Figure 2 (Right) also shows the user-side carbon footprint of downloading datasets from Hugging Face to be comparable to the annual carbon footprint of about 170,000 people (right vertical axis).
@@ 4.2. Social Costs
We analyse the labour-related social costs due to hyper-datafication based on the responses from 134 data workers in Kenya. Respondents report engaging in a wide range of tasks, including data labelling (120/134), content moderation (67/134), data cleaning (52/134), and verification of user preferences (33/134), with many performing multiple task types concurrently. A majority of the respondents (79/134) report that they currently work or have previously worked directly for a major technology company, including OpenAI (51/134), Meta (35/134), Google (24/134), Microsoft (13/134), and Amazon (10/134).
@@
Exposure to graphic content emerges as a central social cost. A majority of respondents (89/134) report some level of exposure, with 60 reporting daily exposure and 25 reporting exposure every other day. This exposure is especially common among workers engaged in content moderation, which is increasingly rebranded as “Trust and Safety”.
Contrary to common claims, our data suggests that higher exposure does not correspond to higher pay. Figure 4 (Right) shows no consistent association between exposure level and salary. Conditioning on working hours and experience yields the same result.
@@
Representation in Datasets. While hyper-datafication increases the volume of data available, this does not necessarily translate into better representation of the world. We use the Hugging Face datasets to investigate this further.
Figure 1 (Left and Centre) shows that text data is a large fraction of the diverse data modalities on Hugging Face. Using these text datasets, we first examine the representation of global languages in datasets hosted on the Hugging Face Hub. Among 57,484 datasets with language annotations, we identify 7,934 distinct languages. Most languages (96%) appear only within multilingual datasets and typically in a small number of repositories. The remaining 4% correspond to 306 languages, of which 97 occur only in a single dataset. This indicates that many languages exist in the margins of the dataset ecosystem rather than as sustained data sources.
@@ 5. Discussions
Lack of Transparency. The environmental costs of hyper-datafication presented in Section 4.1 are based on a snapshot of 550,000 datasets on Hugging Face Hub. Figure 1 shows how the scale, size, and diversity of datasets are growing in recent years. Figure 2 captures the increase in storage-related operational energy consumption and carbon footprint. However, these estimations do not provide a comprehensive overview of the environmental impact of hyper-datafication. Factors such as the embodied emissions, fresh water to cool data centres, raw materials and minerals used in hardware manufacturing, and the e-waste generated due to hardware disposal are not considered (Falk et al., 2025; Selvan, 2025). These broader environmental impacts are difficult to estimate, as acknowledged by works attempting to do this for AI model development (Luccioni et al., 2023; Wright et al., 2023). Lack of transparency and openness in reporting environmental costs by vendors, manufacturers, and actors at all stages of the data lifecycle makes comprehensive environmental cost reporting a challenge.
In addition to the environmental costs, social costs of data, such as labour, are seldom reported and discussed. Results from the data worker questionnaire in Section 4.2 highlight these trends where large corporations employ data workers for diverse types of data work. Although longer hours and greater experience are associated with higher earnings, as shown in Figure 4, salaries remain low relative to the Kenyan national average of USD 540 per month (CEIC Data, 2024). On the one hand, data work creates income opportunities in regions with limited formal employment. [^8] For example, the youth (15-34 years old) in Kenya face an unemployment rate of 67% (Federation of Kenya Employers, ). On the other hand, these same structural conditions enable exploitation (Shitawa, 2024). The availability of surplus labour, combined with limited regulatory oversight, allows data work to be organised through precarious contracts, low wages, and intensive performance monitoring. Economic necessity weakens workers’ bargaining power, while task fragmentation and platform-based subcontracting obscure accountability. Transparent reporting of the costs pertaining to data work can give due credit to the data workers who remain unseen in the discourse surrounding frontier AI. Additionally, the due credit and visibility can help them secure better work conditions and bargaining rights.
@@
However, a frugal alternative exists. Techniques such as representative subset selection and coresets exploit the well-established fact that not all data is equally informative (Katharopoulos and Fleuret, 2018). This allows models to achieve comparable performance with orders of magnitude less data and computation (Killamsetty et al., 2021). However, under current economic incentives, there is little motivation for companies to adopt such approaches. Data accumulation increases not only technical capacity but also strategic control and market valuation, encouraging data hoarding even when its utility is low.
For this reason, data frugality cannot be left to voluntary optimisation. It must be enforced as a structural constraint on hyper-datafication. A corporate data tax could internalise the social and environmental costs of data accumulation, correcting a market failure in which companies benefit from scale while communities and workers bear the burdens. Thus, the tax could discourage wasteful hoarding while generating revenue to compensate data generators and data workers. Frugality, in this sense, is not merely an efficiency principle but a corrective action to the extractive political economy of data (Pasquinelli, 2023).
@@
In addition to the Data PROOFS recommendations, collective action is indispensable to counter the consequences of hyper-datafication and to improve the sustainability of data in frontier AI. Community resistance is growing; in the US, more than 140 activist groups across 24 states are mobilising against new data centre construction and expansion (Data Center Watch, 2025). Over the past two years, these groups have collectively stalled or cancelled proposed projects worth USD 162 billion. Citizen-led protests against hyper-scale data centres are appearing in Mexico, Spain, the Netherlands, India, and other parts of the world (Paul Mozur and Mega, 2025). This trend highlights how the environmental, social, and economic costs of data centre development are becoming increasingly visible, and how local communities are contesting the uneven distribution of benefits and harms associated with hyper-scale AI development.
@@
Data work is carried out globally. However, our questionnaire was only sent to data workers in Kenya. We have refrained from making generalised claims and tried to limit the scope of labour discussions to the conditions in Kenya.
Likewise, some of the economic costs are primarily US-focused. This is again due to the availability of data sources. Furthermore, it reflects the dominance of a few regions, such as the US, which host the majority of global infrastructure for frontier AI development. Additional clarifications about the limitations of using these data sources are provided in Appendix D.4.
@@ 6. Conclusions
The push for hyper-datafication is tied to the objective of modelling reality from data. Terms like digital twins are used widely in this context, aiming to mimic consumer behaviour (Toubia et al., 2025), social interactions (Hendler and Mulvehill, 2016), virtual human twin (European Commission, 2023), planetary climate (Mardani et al., 2023; Doblas-Reyes et al., 2025), and even human life events like death (Savcisens et al., 2024) from large-scale data. We are progressing from an era in which models relied mainly on passively accumulated data to one in which the world is actively reshaped to generate more data. Sensors, platforms, and infrastructures increasingly optimise for continuous data extraction. While hyper-datafication increases the amount of available data, it does not necessarily yield informative and useful data. Furthermore, hyper-datafication does not alleviate existing problems with data-driven models but can aggravate them.
@@ Acknowledgements.
SW and RS acknowledge funding from the Independent Research Fund Denmark (DFF) under grant agreement No. 4307-00143B. RS and JK acknowledge funding from the European Union’s Horizon Europe Research and Innovation Action programme (HORIZON-CL4-2021-HUMAN-01) under grant agreement No. 101070408, project SustainML (Application Aware, Life-Cycle Oriented Model-Hardware Co-Design Framework for Sustainable, Energy Efficient ML Systems). RS further acknowledges funding from the European Union’s Horizon Europe Research and Innovation Action programme under grant agreements No. 101070284 and No. 101189771. ED acknowledges funding from the Novo Nordisk Foundation under grant agreement No. NNF24OC0090977. SM acknowledges funding from the Danish Data Science Academy (DDSA) through DDSA Visit Grant No. 2025-5673. The authors thank the data workers who participated in the anonymous questionnaire for their time and contributions, and members of SAINTS Lab for valuable discussions.
@@ A.1. Metadata Coverage
dataset_size Local storage footprint of the dataset (Apache Arrow, bytes) 74.7
@@ Appendix C Questionnaire Summary Statistics
Free-text responses for type of data work were not included in the analysis. These responses were all unique single mentions (e.g. “AI training”, “customer service”, “3D”).
For employment relationships with large technology companies, free-text responses were used to clarify indirect work arrangements (e.g. work conducted via business process outsourcing (BPO) firms). All such responses were mapped to the predefined category None of the above, reflecting indirect rather than direct employment.
@@
Female 57 ¡1 5 ¡50 15 Verification of user preferences 33
@@ D.2. Model for Storage-related Emissions
with $f_{\text{stored}}=0.1$ denoting the assumed share of downloads that remain stored, $T_{\text{user}}=0.25$ years the assumed average retention time, $D_{i}$ the total number of downloads, which we treat as uniformly distributed over the dataset’s lifetime, and $S_{\text{user},i}$ is the local dataset size. These assumptions reflect limited visibility into actual user-side behaviour and should be interpreted as indicative rather than exact. Based on the available attributes in Table 1, 91% of datasets are included for the provider-side estimate and 75% datasets included for the user-side estimate. Emissions are derived using the average US grid intensity of 384 gCO2eq/kWh (Ritchie et al., 2023a), equivalent to roughly 23 kgCO2eq/TB.
@@ D.3. Internet Traffic and Data Curation
D.4. External Data Source Limitations
All external sources are subject to their respective limitations. Financial estimates rely on investigative reporting rather than audited disclosures, and economic relationships between AI developers and cloud providers are complex and evolving. The geographic concentration of data centres captures presence but not size or utilisation. Accordingly, we use these sources to characterise structural dynamics and support qualitative interpretation, rather than to provide exhaustive or definitive quantification.

### v3

@@ 3.3. External Data Sources
We extract global annual investment in data centre infrastructure from the IEA (International Energy Agency, 2025b). To characterise the geographic concentration and expansion of data centres, we use records from the Data Center Map (19). We assessed electricity demand associated with data centres using historical and projected estimates from the IEA (International Energy Agency, 2024a) and focus on electricity consumption as a primary indicator of environmental impact associated with data centre expansion. We source the carbon intensity for electricity generation from Our World in Data (Ember and Energy Institute, 2025).
@@ 4.1. Environmental Costs
Figure 2. Left: Estimated provider-side storage energy (GWh). Right: Estimated user-side storage energy (TWh), assuming that 10% of downloads result in three months of local storage. To provide scale, cumulative energy is expressed relative to the global average annual per-capita carbon footprint of 4.73 tCO2eq (Ritchie et al., 2023b). For the provider-side, we use the US grid intensity of 384 gCO2eq/kWh (Ember and Energy Institute, 2025) reflecting that nearly all datasets are hosted in the US, while for the user-side we use the global average of 473 gCO2eq/kWh (Ember and Energy Institute, 2025) due to the lack of geographic information on downloads. Provider-side storage thus corresponds to the annual carbon footprint of approximately 80 people and user-side storage to that of roughly 217,000 people.
@@
Figure 2 reports estimated provider-side (Left) and user-side storage (Right) energy consumption associated with Hugging Face datasets. These estimates cover 91% of datasets for the provider-side estimate and 75% for the user-side estimate and is based on assumptions and calculations provided in Appendix D.2. Provider-side storage energy consumption remains modest in absolute terms but increases steadily as datasets accumulate, having reached approximately 1 GWh. User-side costs are approximately three orders of magnitude larger (>>2 TWh), driven by the repeated number of downloads across a large user base. Here, we have assumed that 10% of downloads result in three months of persistent storage per user. These choices are conservative and our estimate should therefore be interpreted as a lower bound.
@@
Carbon intensity varies across electricity grids due to the differences in the energy mix. Storing all datasets hosted on the Hugging Face Hub in the EU rather than the US would reduce the associated carbon footprint by approximately 38%, due to the lower average carbon intensity in the EU (237 gCO2eq/kWh) compared with the US (384 gCO2eq/kWh) in 2024 (Ember and Energy Institute, 2025). Depending on the location of data centre infrastructure, storage-related carbon footprints can differ by up to a factor of 30 across regions [^7] This reflects contrasts between high- and low-carbon electricity grids, such as Kosovo (959 gCO2eq/kWh) and Norway (31 gCO2eq/kWh) in 2024 (Ember and Energy Institute, 2025)..
The user-side carbon footprint of storing local copies of datasets from the Hugging Face Hub is found to be comparable to the annual carbon footprint of about 217,000 people.
@@ 4.2. Social Costs
We analyse labour-related social costs due to hyper-datafication based on the responses from 134 data workers in Kenya. Respondents report engaging in a wide range of tasks, including data labelling (120/134), content moderation (67/134), data cleaning (52/134), and verification of user preferences (33/134), with many performing multiple task types concurrently. A majority (79/134) report that they currently work or have previously worked directly for a major technology company, including OpenAI (51/134), Meta (35/134), Google (24/134), Microsoft (13/134), and Amazon (10/134).
@@
Exposure to graphic content emerges as a central social cost. A majority of respondents (89/134) report some level of exposure, with 60 reporting daily exposure and 25 reporting exposure every other day. This exposure is especially common among workers engaged in content moderation, which is increasingly rebranded as “Trust and Safety”. Contrary to common claims, our data suggests that higher exposure does not correspond to higher pay. Figure 4 (Right) shows no consistent association between exposure level and salary. Conditioning on working hours and experience yields the same result.
@@
Representation in Datasets. While hyper-datafication increases the volume of data available, this does not necessarily translate into better representation of the world. We use the datasets stored on Hugging Face to investigate this further.
Figure 1 (Left and Centre) shows that text data is a large fraction of the diverse data modalities on Hugging Face. Using these text datasets, we first examine the representation of global languages. Among 57,484 datasets with language annotations, we identify 7,934 distinct languages. Most languages (96%) appear only within multilingual datasets and typically in a small number of repositories. The remaining 4% correspond to 306 languages, of which 97 occur only in a single dataset. This indicates that many languages exist in the margins of the dataset ecosystem rather than as sustained data sources.
@@ 5. Discussions
Lack of Transparency. The environmental costs of hyper-datafication presented in Section 4.1 are based on a snapshot of 550,000 datasets on the Hugging Face Hub. Figure 1 shows how the scale, size, and diversity of datasets are growing in recent years. Figure 2 captures the increase in storage-related operational energy consumption and carbon footprint. However, these estimations do not provide a comprehensive overview of the environmental impact of hyper-datafication. Factors such as the embodied emissions, fresh water to cool data centres, raw materials and minerals used in hardware manufacturing, and the e-waste generated due to hardware disposal are not considered (Falk et al., 2025; Selvan, 2025). These broader environmental impacts are difficult to estimate, as acknowledged by works attempting to do this for AI model development (Luccioni et al., 2023; Wright et al., 2023). Lack of transparency and openness in reporting environmental costs by vendors, manufacturers, and actors at all stages of the data lifecycle makes comprehensive environmental cost reporting a challenge.
In addition to the environmental costs, social costs of data, such as labour, are seldom reported and discussed. Results from the data worker questionnaire in Section 4.2 highlight these trends where large corporations employ data workers for diverse types of data work. Although longer hours and greater experience are associated with higher earnings, as shown in Figure 4, salaries remain low relative to the Kenyan national average of USD 540 per month (CEIC Data, 2024). On the one hand, data work creates income opportunities in regions with limited formal employment. [^8] For example, the youth (15-34 years old) in Kenya face an unemployment rate of 67% (Federation of Kenya Employers, ). On the other hand, these same structural conditions enable exploitation (Shitawa, 2024). The availability of surplus labour, combined with limited regulatory oversight, allows data work to be organised through precarious contracts, low wages, and intensive performance monitoring. Economic necessity weakens workers’ bargaining power, while task fragmentation and platform-based subcontracting obscure accountability. Transparent reporting of the costs pertaining to data work can give due credit to the data workers who remain unseen in the discourse surrounding frontier AI. Additionally, the credit and visibility can help them secure better work conditions and bargaining rights.
Although enforcing such reporting is not straightforward, it could be encouraged through incentive structures that offer companies strategic recognition for disclosing data-related environmental and social impacts. For example, certification of sustainable data practices could reward companies for comprehensive reporting. This could frame transparency as a competitive advantage rather than merely a regulatory burden.
@@
However, a frugal alternative exists, which builds on the well-established fact that not all data is equally informative (Katharopoulos and Fleuret, 2018). Techniques such as representative subset selection and coresets identify a small, representative, and potentially reweighted subset of data that approximately preserves the learning dynamics of a full dataset. This allows models to achieve comparable performance with substantially less data and computation (Killamsetty et al., 2021; Sorscher et al., 2022; Muennighoff et al., 2023) and reduced environmental cost (Wilson et al., 2026). Despite these benefits, these approaches often require one-time upfront construction effort, which is amortised when the resulting subset is sufficiently small or used multiple times, for example, to tune model hyperparameters. Nevertheless, under current economic incentives, there is little motivation for companies to adopt such approaches. Data accumulation increases not only technical capacity but also strategic control and market valuation, encouraging data hoarding even when its utility is low. Finally, companies may simply retain data due to a “fear of missing out” on valuable insights or future utility (Gormley and Gormley, 2012; Zuboff, 2023).
For this reason, data frugality cannot be left to voluntary optimisation. It must be enforced as a structural constraint on hyper-datafication. A corporate data tax could internalise the social and environmental costs of data accumulation, correcting a market failure in which companies benefit from scale while communities and workers bear the burdens. Thus, the tax could discourage wasteful hoarding while generating revenue to compensate data generators and data workers, and increase transparency around data ownership and use. A related proposal by Irwin et al. (2026) suggests a data “token tax” on AI inference to capture value where AI is used rather than where the infrastructure is hosted. Tax-based instruments are, however, most often jurisdiction-specific and risk displacing activity to data tax havens. Effective implementation therefore requires frameworks with extraterritorial reach and binding legal force, such as the EU AI Act, which could serve as a basis for introducing data cost-related accountability mechanisms. Frugality, in this sense, is not merely an efficiency principle but a corrective action, which must be backed by enforceable policy to meaningfully change the extractive political economy of data (Pasquinelli, 2023).
@@
In addition to the Data PROOFS recommendations, collective action is indispensable to counter the consequences of hyper-datafication and to improve the sustainability of data in frontier AI. Community resistance is growing; in the US, more than 140 activist groups across 24 states are mobilising against new data centre construction and expansion (Data Center Watch, 2025). Over the past two years, these groups have collectively stalled or cancelled proposed projects worth USD 162 billion. Citizen-led protests against hyper-scale data centres are appearing in Mexico, Spain, the Netherlands, India, and other parts of the world (Mozur et al., 2025). This trend highlights how the environmental, social, and economic costs of data centre development are becoming increasingly visible, and how local communities are contesting the uneven distribution of benefits and harms associated with hyper-scale AI development.
@@
Data work is carried out globally, but our questionnaire was limited to workers in Kenya. We have refrained from making generalised claims and tried to limit the scope of labour discussions to the conditions in Kenya.
Likewise, some of the economic costs are primarily US-focused. This is again due to the availability of data sources. Furthermore, it reflects the dominance of a few regions, such as the US, which host the majority of global infrastructure for frontier AI development. Finally, while the geographic concentration of data centres captures their presence, it does not reflect their purpose, scale or utilisation.
@@ 6. Conclusions
The push for hyper-datafication is tied to the objective of modelling reality from data. In this pursuit, we are progressing from an era in which models relied mainly on passively accumulated data to one in which the world is actively reshaped to generate more data. Sensors, platforms, and infrastructures increasingly optimise for continuous data extraction. While hyper-datafication increases the amount of available data, it does not necessarily yield informative and useful data. Furthermore, hyper-datafication does not alleviate existing problems with data-driven models but can aggravate them.
@@ Acknowledgements.
SW and RS acknowledge funding from the Independent Research Fund Denmark (DFF) under grant agreement No. 4307-00143B. RS and JK acknowledge funding from the European Union’s Horizon Europe Research and Innovation Action programme (HORIZON-CL4-2021-HUMAN-01) under grant agreement No. 101070408, project SustainML (Application Aware, Life-Cycle Oriented Model-Hardware Co-Design Framework for Sustainable, Energy Efficient ML Systems). RS further acknowledges funding from the European Union’s Horizon Europe Research and Innovation Action programme under grant agreements No. 101070284 and No. 101189771. ED acknowledges funding from the Novo Nordisk Foundation under grant agreement No. NNF24OC0090977. SM acknowledges funding from the Danish Data Science Academy (DDSA) through DDSA Visit Grant No. 2025-5673. The authors thank the data workers who participated in the anonymous questionnaire for their time and contributions, and members of SAINTS Lab for valuable discussions. The authors also thank the anonymous reviewers and the area chair at FAccT for their valuable feedback.
Author Contributions
Contributions follow the CRediT taxonomy [^11] https://credit.niso.org/. SW: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Project administration, Resources, Software, Validation, Visualization, Writing – original draft, Writing – review & editing. SM: Methodology, Validation, Writing – original draft, Writing – review & editing. MO: Investigation, Resources. ED: Conceptualization, Supervision, Validation, Writing – original draft, Writing – review & editing. JK: Conceptualization, Data curation, Methodology, Resources, Validation, Writing – original draft, Writing – review & editing. RS: Conceptualization, Data curation, Investigation, Methodology, Project administration, Resources, Supervision, Validation, Visualization, Writing – original draft, Writing – review & editing.
@@ A.1. Metadata Coverage
dataset_size Local storage footprint of the dataset (bytes) 74.7
@@ Appendix C Questionnaire Summary Statistics
Free-text responses for type of data work were not included in the analysis. These responses were all unique single mentions (e.g., “AI training”, “customer service”, “3D”).
For employment relationships with large technology companies, free-text responses were used to clarify indirect work arrangements (e.g., work conducted via business process outsourcing (BPO) firms). All such responses were mapped to the predefined category None of the above, reflecting indirect rather than direct employment.
@@
Female 57 ¡1 5 ¡50 15 Labelling data 120
@@ D.2. Model for Storage-related Emissions
with $f_{\text{stored}}=0.1$ denoting the assumed share of downloads that remain stored, $T_{\text{user}}=0.25$ years the assumed average retention time, $D_{i}$ the total number of downloads, which we treat as uniformly distributed over the dataset’s lifetime, and $S_{\text{user},i}$ is the local dataset size. These assumptions reflect limited visibility into actual user-side behaviour and should be interpreted as indicative rather than exact. Based on the available attributes in Table 1, 91% of datasets are included for the provider-side estimate and 75% datasets included for the user-side estimate.
For provider-side storage, emissions are derived using the average US grid intensity of 384 gCO2eq/kWh (Ritchie et al., 2023a), as 552,713 of the datasets with region information (99.8%) are hosted in the US, compared to only 316 in the EU. This corresponds to approximately 23.0 kgCO2eq/TB. For user-side storage, we apply the global average grid intensity of 473 gCO2eq/kWh (Ritchie et al., 2023a), given the lack of geographic information on dataset downloads, yielding an estimate of roughly 28.4 kgCO2eq/TB.